OpenStudy (elson):

Write the complete balanced equation for the following reaction: C5H9O + O2 yields CO2 + H2O please help, i'm lost :/

OpenStudy (hlilly2413):

Ok, so do you know to start by making sure there are equal amounts of C, H, and O on each side of the equation? On the reactant side (left side) you start with 5 atoms of C, 9 atoms of H, and 3 atoms of O. On the product side (right side) you start with 1 atom of C, 2 atoms of H, and 3 atoms of O. The only equal atom is C. However, you need to get all atoms equal. So, I would start with C by placing a 5 in front of the CO2 on the product side. Then, you have 5 carbon atoms on each side, but your # of O atoms changes. Leave that for now and go to H. Balance the hydrogen, etc. etc. You may hit a wall at a certain point, but perhaps you won't. Try it and post how far you get. Hope this has helped a little.

OpenStudy (hlilly2413):

Opps, the only equal atom is O to start, not C. Sorry for the typo.**

OpenStudy (welshfella):

start with 5CO2 on right side C5H9O + O2 = 5CO2 + H2O there are 9 H's on left so make if even by multiply by first compound by 2 and write 9H2O 2 C5H9O + O2 = 10CO2 + 9H2O Note I had to makw it 10CO2

OpenStudy (welshfella):

now there are 4 O's on the left and 29 on the right...

OpenStudy (welshfella):

OK make it 18 H2O and 4C5H9O 4C5H9O + O2 = 20CO2 + 18H2O so on the left we have 6 O's and 58 O's on the right so make it 27O2:- 4C5H9O + 27O2 ---> 20CO2 + 18H2O does that balance?

OpenStudy (welshfella):

Yep

OpenStudy (elson):

thanks everyone!!:)

OpenStudy (hlilly2413):

Ty for the medal, I didn't want to do it for you. Glad you were able to get your answer!